# Unlearning as multi-task optimization: A normalized gradient difference approach with an adaptive learning rate

### v1

@@ 5.1 Setup
Task of Fictitious Unlearning (TOFU) (Maini et al., 2024). TOFU consists of 20 question-answer pairs based on fictitious author biographies generated by GPT-4 (Achiam et al., 2023). In our experiments, we use the forget10 (10% of the full training set) as the forgetting set and retain90 (90% of the full training set) as the retaining set.
MUSE-NEWS (Shi et al., 2024). This dataset consists of BBC news articles (Li et al., 2023b) published since August 2023. We use its train split to finetune a target model, and then the raw set, which includes both the forgetting and retaining data, for the target model unlearning. Finally, the verbmem and knowmem splits are used to evaluate the unlearned model’s performance.
Unlearning methods We compare NGDiff with 4 baselines. The first baseline method is the target model without any unlearning, while the remaining three are the state-of-the-art unlearning methods.
@@
Gradient Difference (GDiff) (Liu et al., 2022). GDiff (see Sec. 3.2) applies static linear scalarization with $c=0.5$ in MTO. For a thorough comparison, we also include the extended GDiff method, with $c=0.1$ or $c=0.9$.
@@
Negative Preference Optimization (NPO) (Zhang et al., 2024b). NPO uses preference optimization (Ouyang et al., 2022) with the loss: $L_{\text{NPO},\beta}(\bm{\theta})=$
@@
We test multiple LLMs: LLaMA2-7B (Touvron et al., 2023), Phi-1.5 (Li et al., 2023a), Falcon-1B (Penedo et al., 2023), GPT2-XL (Radford et al., 2019) and Mistral-7B (Jiang et al., 2023). They are pre-trained and then fine-tuned on datasets in Section 5.1, with AdamW optimizer and are carefully tuned (Appendix F).
@@ 5.2 Evaluation Metrics
Following the existing work (Shi et al., 2024), we evaluate the unlearning performance based on model’s output quality. We expect a good performance should satisfy the following requirements:
@@ 6 Conclusion and Discussion
Ethical Consideration
We acknowledge that our work is aligned with the ACL Code of the Ethics  [^8] https://www.aclweb.org/portal/content/acl-code-ethics and will not raise ethical concerns. We do not use sensitive datasets/models that may cause any potential issues/risks.
@@ Appendix B Comparing NGDiff with GradNorm
6:     Construct unlearning gradient $\bm{g}_{\textup{NGDiff}}=\bm{g}_{\textup{R}}/||\bm{g}_{\textup{R}}||-\bm{g}_{\textup{F}}/||\bm{g}_{\textup{F}}||$
@@
We compare the GradNorm algorithm Chen et al. (2018b) with our proposed method, NGDiff. We highlight some steps of GradNorm in red to indicate the differences than NGDiff:
  • NGDiff sets the scalaring coefficient as $1/||\bm{g}_{\textup{R}}||$ and $1/||\bm{g}_{\textup{F}}||$, while GradNorm uses gradient descent to learn these coefficients as $w_{\textup{R}}$ and $w_{\textup{F}}$.
@@ C.1 Brief introduction of GeN
GeN (Bu and Xu, 2024) is a method that sets the learning rate for any given gradient $\bm{d}$ as
@@ Appendix D Computer Vision Experiments
To demonstrate the effectiveness of unlearning across other modalities, we also evaluate our method on the image classification task. Specifically, we choose the CIFAR-10 and CIFAR-100 dataset Krizhevsky et al. (2009) and train a ResNet-50 He et al. (2015) model from scratch. For the CIFAR-10 dataset, we sample 500 images from the class $dog$ as the forgetting data, and use images from the remaining 9 classes as the retaining data. For the CIFAR-100 dataset, we sample 500 images from the class $bed$ as the forgetting data, and use images from the remaining 99 classes as the retaining data. After training, the initial forget data accuracy is $0.926$, and the retain data accuracy is $0.956$ on the CIFAR-10 dataset. The initial forget data accuracy is $0.745$, and the retain data accuracy is $0.750$ on the CIFAR-100 dataset. Then we apply different unlearning methods to the trained models. As shown in Table 4, all methods successfully reduce the forget accuracy to 00. However, the retaining accuracy of NGDiff remains the highest, which shows its effectiveness in preserving the model utility in image classification tasks.
@@ Appendix E Other Related Works
Machine unlearning is oftentimes viewed as a continual learning approach, that removes specific data points after a model has been trained to memorize them. Such removal is light-weighted in contrast to re-training, especially when the forgetting set is much smaller than the retaining. In addition to the methods already introduced in Section 3.2 (namely GA, GDiff and NPO), other methods include SISA Bourtoule et al. (2021), influence functions Ullah et al. (2021), differential privacy Gupta et al. (2021) and so on. However, these methods could be difficult to scale on large models and large datasets due to the algorithmic complexity. To our best knowledge, this is the first work that formulate the unlearning problem as a two-task problem, which can be solved by a number of well-known MTO methods.
@@
MTO is a paradigm where one model is trained to perform multiple tasks simultaneously, so as to significantly improve the efficiency in contrast to training multiple models, one for each task. The key challenge of MTO is the performance trade-off among tasks, where the multi-task model is worse than single-task model if trained on each task separately. Therefore, the core idea is to balance different tasks by modifying the per-task gradients, e.g. with normalization (LossNorm and NGDiff), PCGrad Yu et al. (2020), RLW Lin et al. (2021), IMTL Liu et al. (2021), MGDA Désidéri (2012), CAGrad Liu et al. (2024a), GradVaccine Wang et al. (2020), GradDrop Chen et al. (2020), RotoGrad Javaloy and Valera (2022), etc.
@@
On one hand, GeN Bu and Xu (2024) leverages the Taylor expansion and convex-like landscape of deep learning, which is applicable for the general purpose, even if the gradient is modified like in the unlearning.
On the other hand, methods like D-adaptation Defazio and Mishchenko (2023), Prodigy Mishchenko and Defazio (2024), DoG Ivgi et al. (2023), DoWG Khaled et al. (2024) are based on the convex and $G$-Lipschitz conditions: $L(\bar{\theta}_{T})-L(\bm{\theta}_{*})\leq\frac{|\theta_{0}-\theta_{*}|^{2}}{2\eta T}+\frac{\eta G^{2}}{2}$ where $\theta_{*}$ is the unknown minimizer of $L$ and $\bar{\theta}_{T}$ is an averaging scheme of $\{\theta_{0},...,\theta_{T}\}$. With the same theoretical foundation, these methods propose different ways to approximate the initial-to-final distance $|\theta_{0}-\theta_{*}|$. There are two main issues to apply these methods on the unlearning. Firstly, the assumption of $G$-Lipschitz is hard to verify and the minimizer $\theta_{*}$ is not well-defined in multi-objective (see our discussion on Pareto optimality under Lemma 2). Secondly, the optimal learning rate $\frac{|\theta_{0}-\theta_{*}|}{G\sqrt{T}}$ is defined in a manner to minimize the loss, whereas MTO methods operate on the gradient level. Hence MTO is incompatible to such parameter-free methods given that we cannot derive a corresponding loss (e.g. there exists no $L_{\textup{NGDiff}}$ such that $\frac{\partial L_{\textup{NGDiff}}}{\partial\theta}=\bm{g}_{\textup{NGDiff}}$).
@@ F.1 Datasets
  • Task of Fictitious Unlearning (TOFU) Maini et al. (2024). This dataset consists of question-answer pairs based on fictitious author biographies generated by GPT-4 Achiam et al. (2023). Initially, predefined attributes, such as birthplace, gender, and writing genre, are assigned to 200 distinct authors. GPT-4 is then prompted to generate detailed information about each author. Following the synthesized data, 20 question-answer pairs are created for each fictitious author. The dataset is then divided into distinct datasets: the retaining set and the forgetting set. In our experiments, we use the forget10 and retain90 split, which excludes 10% of the original dataset.
  • MUSE-NEWS Shi et al. (2024). This dataset consists of BBC news articles Li et al. (2023b) from August 2023. It includes seven subsets of news data: raw, verbmem, knowmem, privleak, scal, sust, and train. We utilize the train split to finetune a target model, and then the raw set, which includes both the forget and retain data, for the target model unlearning. Then, we use verbmem, knowmem split to evaluate the unlearned model’s performance.
@@ F.2 Evaluation Metrics
Following the existing work Shi et al. (2024), we evaluate the unlearning performance based on the quality of outputs from the model after unlearning. We expect a good performance should satisfy the following requirements:
@@
$\begin{split}&\textit{Verbmem}(f,F)\\ &=\frac{1}{||F||}\Sigma_{x}\>\textsc{ROUGE-L}(f(x_{[:l]}),x_{[l+1:]})\end{split}$ (9)
@@
$\textit{Knowmem}(f,F)=\frac{1}{||F||}\Sigma_{x}\>\textsc{ROUGE-L}(f(q),a)$ (10)
@@
$\textit{Utility}(f,R)=\frac{1}{||R||}\Sigma_{x}\>\textsc{ROUGE-L}(f(q),a)$ (11)
@@ Appendix G Proofs
Lemma 2 (restated from Xin et al. (2022)).
@@
$\begin{split}\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}&=\bm{g}_{\textup{R}}^{\top}\left(\frac{\bm{g}_{\textup{R}}}{||\bm{g}_{\textup{R}}||}-\frac{\bm{g}_{\textup{F}}}{||\bm{g}_{\textup{F}}||}\right)\\ &=||\bm{g}_{\textup{R}}||-\frac{\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{F}}}{||\bm{g}_{\textup{F}}||}\geq||\bm{g}_{\textup{R}}||-\frac{||\bm{g}_{\textup{R}}||||\bm{g}_{\textup{F}}||}{||\bm{g}_{\textup{F}}||}\\ &=0\end{split}$
@@
$L_{\textup{R}}(\bm{\theta}_{t+1})-L_{\textup{R}}(\bm{\theta}_{t})$
$=-\eta\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}+\frac{\eta^{2}}{2}\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}_{\textup{R}}\bm{g}_{\textup{NGDiff}}+o(\eta^{2})$ (12)
@@
$\begin{split}&-\eta\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}+\frac{\eta^{2}}{2}\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}_{\textup{R}}\bm{g}_{\textup{NGDiff}}<0\\ &\Longleftrightarrow 0<\eta<\frac{2\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}}{\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}_{\textup{R}}\bm{g}_{\textup{NGDiff}}}\end{split}$

### v2

@@ 5.1 Setup
Task of Fictitious Unlearning (TOFU) [36]. TOFU consists of 20 question-answer pairs based on fictitious author biographies generated by GPT-4 [1]. In our experiments, we use the forget10 (10% of the full training set) as the forgetting set and retain90 (90% of the full training set) as the retaining set.
MUSE-NEWS [43]. This dataset consists of BBC news articles [29] published since August 2023. We use its train split to finetune a target model, and then the raw set, which includes both the forgetting and retaining data, for the target model unlearning. Finally, the verbmem and knowmem splits are used to evaluate the unlearned model’s performance.
Unlearning methods We compare NGDiff with 4 baselines. The first baseline method is the target model without any unlearning, while the remaining three are popular unlearning methods.
@@
Gradient Difference (GDiff) [31]. GDiff (see Sec. 3.2) applies static linear scalarization with $c=0.5$ in MTO. For a thorough comparison, we also include the extended GDiff method, with $c=0.1$ or $c=0.9$.
@@
Negative Preference Optimization (NPO) [55]. NPO uses preference optimization [38] with the loss: $L_{\text{NPO},\beta}(\bm{\theta})=$
@@
We test multiple LLMs: LLaMA2-7B [46], Phi-1.5 [28], Falcon-1B [40], GPT2-XL [41] and Mistral-7B [24]. They are pre-trained and then fine-tuned on datasets in Section 5.1, with AdamW optimizer and are carefully tuned (Appendix F).
@@ 5.2 Evaluation Metrics
Following the existing work [43], we evaluate the unlearning performance based on model’s output quality. We expect a good performance should satisfy the following requirements:
@@ Appendix A Preliminary evidence
Figure 6: ROUGE scores and loss values during unlearning with vanilla GDiff (equally weighted), under different learning rates to which the unlearning performance is highly sensitive.
@@ Appendix B Comparing NGDiff with GradNorm
6:    Construct unlearning gradient $\bm{g}_{\textup{NGDiff}}=\bm{g}_{\textup{R}}/\|\bm{g}_{\textup{R}}\|-\bm{g}_{\textup{F}}/\|\bm{g}_{\textup{F}}\|$
@@
We compare the GradNorm algorithm [11] with our proposed method, NGDiff. We highlight some steps of GradNorm in red to indicate the differences than NGDiff:
  • NGDiff sets the scalaring coefficient as $1/\|\bm{g}_{\textup{R}}\|$ and $1/\|\bm{g}_{\textup{F}}\|$, while GradNorm uses gradient descent to learn these coefficients as $w_{\textup{R}}$ and $w_{\textup{F}}$.
@@ C.1 Brief introduction of GeN
GeN [4] is a method that sets the learning rate for any given gradient $\bm{d}$ as
@@ Appendix D Computer Vision Experiments
To demonstrate the effectiveness of unlearning across other modalities, we also evaluate our method on the image classification task. Specifically, we choose the CIFAR-10 and CIFAR-100 dataset [26] and train a ResNet-50 [20] model from scratch. For the CIFAR-10 dataset, we sample 500 images from the class $dog$ as the forgetting data, and use images from the remaining 9 classes as the retaining data. For the CIFAR-100 dataset, we sample 500 images from the class $bed$ as the forgetting data, and use images from the remaining 99 classes as the retaining data. After training, the initial forget data accuracy is $0.926$, and the retain data accuracy is $0.956$ on the CIFAR-10 dataset. The initial forget data accuracy is $0.745$, and the retain data accuracy is $0.750$ on the CIFAR-100 dataset. Then we apply different unlearning methods to the trained models. As shown in Table 4, all methods successfully reduce the forget accuracy to 00. However, the retaining accuracy of NGDiff remains the highest, which shows its effectiveness in preserving the model utility in image classification tasks.
@@ Appendix E Other Related Works
Machine unlearning is oftentimes viewed as a continual learning approach, that removes specific data points after a model has been trained to memorize them. Such removal is light-weighted in contrast to re-training, especially when the forgetting set is much smaller than the retaining. In addition to the methods already introduced in Section 3.2 (namely GA, GDiff and NPO), other methods include SISA [2], influence functions [47], differential privacy [18] and so on. However, these methods could be difficult to scale on large models and large datasets due to the algorithmic complexity. To our best knowledge, this is the first work that formulate the unlearning problem as a two-task problem, which can be solved by a number of well-known MTO methods.
@@
MTO is a paradigm where one model is trained to perform multiple tasks simultaneously, so as to significantly improve the efficiency in contrast to training multiple models, one for each task. The key challenge of MTO is the performance trade-off among tasks, where the multi-task model is worse than single-task model if trained on each task separately. Therefore, the core idea is to balance different tasks by modifying the per-task gradients, e.g. with normalization (LossNorm and NGDiff), PCGrad [52], RLW [30], IMTL [33], MGDA [16], CAGrad [32], GradVaccine [49], GradDrop [12], RotoGrad [23], etc.
@@
On one hand, GeN [4] leverages the Taylor expansion and convex-like landscape of deep learning, which is applicable for the general purpose, even if the gradient is modified like in the unlearning.
On the other hand, methods like D-adaptation [14], Prodigy [37], DoG [21], DoWG [25] are based on the convex and $G$-Lipschitz conditions: $L(\bar{\theta}_{T})-L(\bm{\theta}_{*})\leq\frac{|\theta_{0}-\theta_{*}|^{2}}{2\eta T}+\frac{\eta G^{2}}{2}$ where $\theta_{*}$ is the unknown minimizer of $L$ and $\bar{\theta}_{T}$ is an averaging scheme of $\{\theta_{0},...,\theta_{T}\}$. With the same theoretical foundation, these methods propose different ways to approximate the initial-to-final distance $|\theta_{0}-\theta_{*}|$. There are two main issues to apply these methods on the unlearning. Firstly, the assumption of $G$-Lipschitz is hard to verify and the minimizer $\theta_{*}$ is not well-defined in multi-objective (see our discussion on Pareto optimality under Lemma 2). Secondly, the optimal learning rate $\frac{|\theta_{0}-\theta_{*}|}{G\sqrt{T}}$ is defined in a manner to minimize the loss, whereas MTO methods operate on the gradient level. Hence MTO is incompatible to such parameter-free methods given that we cannot derive a corresponding loss (e.g. there exists no $L_{\textup{NGDiff}}$ such that $\frac{\partial L_{\textup{NGDiff}}}{\partial\theta}=\bm{g}_{\textup{NGDiff}}$).
@@ F.1 Datasets
  • Task of Fictitious Unlearning (TOFU) [36]. This dataset consists of question-answer pairs based on fictitious author biographies generated by GPT-4 [1]. Initially, predefined attributes, such as birthplace, gender, and writing genre, are assigned to 200 distinct authors. GPT-4 is then prompted to generate detailed information about each author. Following the synthesized data, 20 question-answer pairs are created for each fictitious author. The dataset is then divided into distinct datasets: the retaining set and the forgetting set. In our experiments, we use the forget10 and retain90 split, which excludes 10% of the original dataset.
  • MUSE-NEWS [43]. This dataset consists of BBC news articles [29] from August 2023. It includes seven subsets of news data: raw, verbmem, knowmem, privleak, scal, sust, and train. We utilize the train split to finetune a target model, and then the raw set, which includes both the forget and retain data, for the target model unlearning. Then, we use verbmem, knowmem split to evaluate the unlearned model’s performance.
@@ F.2 Evaluation Metrics
Following the existing work [43], we evaluate the unlearning performance based on the quality of outputs from the model after unlearning. We expect a good performance should satisfy the following requirements:
@@
$\begin{split}&\textit{Verbmem}(f,F)\\ &=\frac{1}{\|F\|}\Sigma_{x}\>\textsc{ROUGE-L}(f(x_{[:l]}),x_{[l+1:]})\end{split}$ (10)
@@
$\textit{Knowmem}(f,F)=\frac{1}{\|F\|}\Sigma_{x}\>\textsc{ROUGE-L}(f(q),a)$ (11)
@@
$\textit{Utility}(f,R)=\frac{1}{\|R\|}\Sigma_{x}\>\textsc{ROUGE-L}(f(q),a)$ (12)
@@ Appendix G Proofs
Lemma 2 (restated from [51]).
@@
$\begin{split}\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}&=\bm{g}_{\textup{R}}^{\top}\left(\frac{\bm{g}_{\textup{R}}}{\|\bm{g}_{\textup{R}}\|}-\frac{\bm{g}_{\textup{F}}}{\|\bm{g}_{\textup{F}}\|}\right)\\ &=\|\bm{g}_{\textup{R}}\|-\frac{\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{F}}}{\|\bm{g}_{\textup{F}}\|}\geq\|\bm{g}_{\textup{R}}\|-\frac{\|\bm{g}_{\textup{R}}\|\|\bm{g}_{\textup{F}}\|}{\|\bm{g}_{\textup{F}}\|}\\ &=0\end{split}$
@@
$L_{\textup{R}}(\bm{\theta}_{t+1})-L_{\textup{R}}(\bm{\theta}_{t})=-\eta\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}+\frac{\eta^{2}}{2}\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}_{\textup{R}}\bm{g}_{\textup{NGDiff}}+o(\eta^{2})$ (13)
@@
$\begin{split}&-\eta\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}+\frac{\eta^{2}}{2}\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}_{\textup{R}}\bm{g}_{\textup{NGDiff}}<0\Longleftrightarrow 0<\eta<\frac{2\bm{g}_{\textup{R}}^{\top}\bm{g}_{\textup{NGDiff}}}{\bm{g}_{\textup{NGDiff}}^{\top}\mathbf{H}_{\textup{R}}\bm{g}_{\textup{NGDiff}}}\end{split}$